A point dipole of dipole moment p is located at the center of a grounded ($shel 0) spherical conducting shell of radius R. a Find the electrostatic potential POT) inside the shell.
Answer #1

As the dipole moment p is placed at the center.

We know that, point dipole at origin will have the potential of

(1/4π\epsilon).(p.cos\theta/r2)

Laplace equation :

lrl+ Bl/rl+1

l= 1 is only relevant term ,( held at center)

V(r,\theta)= (A1r1+ p/4π\epsilon0r2)cos\theta +\sum Aerl-1plcos\theta  [ Ae=0]

V(r=R,\theta)= A1R+ (p/4π\epsilon)(1/R2)=0

V(r,\theta)= p/4π\epsilon0(1/r2-r/R3)cos\theta